# DUDF: Differentiable Unsigned Distance Fields with Hyperbolic Scaling

### v1

@@ 1 Introduction
Surface representation is a fundamental aspect in the field of 3D geometry processing, with explicit methods such as triangle meshes, point clouds, and voxelized representations being traditional choices. Implicit surface representations, on the other hand, have been an integral part of the graphics pipeline for many decades. They encapsulate surfaces as the zero-level set of a function, providing a compact and continuous geometry representation. The novelty in recent years has emerged from parameterizing these implicit functions with Neural Networks (NNs), combining their learning capabilities with the advantages of implicit representations.
@@
We evaluate our framework with experiments that rigorously assess performance on a variety of challenging data sets and compare to competitive benchmarks in the field. Our method overcomes several limitations of previous approaches, which struggled with the non-differentiability of UDFs. Results demonstrate that our approach not only enhances the reconstruction quality of open surfaces but also accelerates the training process. We believe the dual improvement in precision and efficiency enhance the applicability of neural UDFs representations in geometric processing for real-world scenarios.
@@ 2 Related Works
Closed surfaces. Recent years have seen a surge in the use of Neural Implicit Functions (NIFs) for modeling and reconstructing 3D shapes [39]. These representations are usually implemented with Multi-Layer Perceptrons (MLP) or Convolutional Neural Networks (CNNs), but differ in the learning task [21, 4, 29, 5]. On the one hand, some methods [21, 4, 29, 22, 5] learn an indicator function or binary occupancies which are later used to reconstruct the 3D surface. On the other hand, alternative methods focus on estimating the SDF at any point in 3D space [27, 15, 23, 9, 11, 24]. The landscape of NIFs has further expanded with the introduction of novel approaches like implicit moving least-squares surfaces [16], a differentiable Poisson solver [28], a complex Gabor wavelet [32], and a level set alignment loss [20]. Notably, central works in the field [13, 35] approached the problem as finding solutions to the Eikonal equation alongside expressive boundary conditions, building upon the literature on solving PDEs with neural networks [34, 30]. In particular, SIREN [35] achieved remarkable results utilizing periodic activation functions, allowing to successfully control the function’s differential fields. Despite these advancements, a fundamental limitation of current NIF approaches remains in their inability to represent open surfaces, a characteristic often exhibited by real-world objects like scene walls, clothing, or vehicles with inner structures.
@@
Open surfaces. In order to model general non-watertight surfaces, Chibane et al. introduced the idea of learning UDFs, pioneering the handling of open surfaces with neural networks [6]. Following this, several methods have aimed to improve the performance of open-surface representation using neural networks [36, 37, 42, 41, 43, 45]. For instance, GIFS [42] and Neural Vector Fields [41] model the relationships between every two points instead of the relationships between points and surfaces. HSDF separately learns sign and distance fields to handle surfaces with arbitrary topologies[38]. NeuralUDF [18] and NeUDF [17] focused on learning an unsigned distance field by volume rendering for multi-view reconstruction of surfaces with arbitrary topologies. Closer to our work is CAP-UDF [45], which optimized models on raw point clouds by learning to move 3D point queries until reaching the surface with a field consistency constraint. Despite these developments, since UDFs are not differentiable at the zero level set, the gradient field is ill-defined near the isosurface. This leads to difficulties during training, but specially during surface reconstruction, where gradients are often used to determine surface presence between points [14, 46], compute surface normals, and project points onto the isosurface [6] through gradient descent, among others. Recent work by Zhou et al. aimed to solve this issue by introducing constraints at the zero level set in the form of losses leading to a smoother surface [44]. In contrast, our approach formulates an Eikonal problem which is solved by learning a scaled distance function that remains differentiable close to the surface. We demonstrate that our method not only enhances reconstruction quality, addressing the prevailing challenges associated with the non-differentiability of the unsigned distance, but also improves state-of-the-art in terms of training time performance.
@@ 3.1 Mathematical background
Methods addressing closed surfaces approach the learning of SDFs as finding the solution to a system of Partial Differential Equations (PDEs) governed by the homogeneous Eikonal equation with Dirichlet and Neumann boundary conditions. Formally, given a closed surface $\mathcal{S}$ in $\mathcal{C}\subseteq\mathbb{R}^{3}$ they seek to find a continuous function $f$ that satisfies $\left\|\nabla f\right\|=1$, with boundary conditions $f_{\rvert_{\mathcal{S}}}=0$, $\nabla f_{\rvert_{\mathcal{S}}}=\mathbf{n}_{\mathcal{S}}$. Where $\mathbf{n}_{\mathcal{S}}$ denotes the unitary normal field at the surface. Although SDFs are not differentiable at every point and represent only a weak solution to the Eikonal equation, recent research has demonstrated considerable success in solving this problem with continuously differentiable implicit neural representation networks [35, 13, 26, 7]. These architectures use periodic activations, facilitating smoother optimization processes and improved control over the solution’s gradient field. This is achievable because locations where the signed distance lacks differentiability are distant from the isosurface, hence the approximation tends to be good in a close neighborhood of the zero level set. In particular, using these networks enable accurate computation of critical topological properties, including mean and Gaussian curvature [26]. This contrasts with networks featuring piecewise linear activations, such as ReLUs, which have null second-order derivatives [35].
@@ 3.2 Problem statement
Our key insight is to overcome these challenges by redefining the unsigned distance field, particularly through the application of an hyperbolic scaling. In light of this understanding, we propose to learn the parameters $\theta$ of a neural network $f_{\theta}$ with periodic activation functions [35] to approximate the function:
@@
where $\mathit{d}_{\mathcal{S}}$ is the unsigned distance to surface $\mathcal{S}$; and $\alpha$ is a constant value. The function $\mathit{t}_{\mathcal{S}}$ is a differentiable approximation of $\mathit{d}_{\mathcal{S}}$, whose zero level set is the surface $\mathcal{S}$. In Fig. 2 (a), we show the impact of our hyperbolic scaling on the unsigned distance field, mainly characterized by quadratic smoothing in proximity to the isosurface and a linear behavior in the distance. The parameter $\alpha$ controls the distance at which the smoothing occurs.
The scaling enables the application of continuously differentiable implicit neural representation networks to solve an Eikonal equation, while retaining the UDF’s open surface representation capabilities. For this task, we aim to address an heterogeneous Eikonal equation, for which we know $t_{\mathcal{S}}$ (Eq. 1) is a weak solution:
@@
The problem outlined in Eq. 2 does not incorporate the information of the surface normals. This contrasts with prior research focused on closed surfaces, where leveraging such information has led to improved accuracy and enhanced reconstruction quality by further constraining the set of feasible solutions. Our second key insight is that although the gradient of $t_{\mathcal{S}}$ (Eq. 1) becomes null at the isosurface, the direction in which the gradient norm increases most rapidly remains non-null. This direction is indicated by the maximum curvature of $f$, defined as the eigenvector associated with the largest eigenvalue of the Hessian matrix. Note that in this context, the maximum curvature does not refer to the surface’s curvature, but rather to the curvature of the hyperbolic scaled unsigned distance field $f$, where the function’s rate of change varies most rapidly. Formally, $\mathbf{H}_{t_{\mathcal{S}}}$ is the Hessian matrix of $t_{\mathcal{S}}$ which has 3 real eigenvalues $\left|\lambda_{1}\right|\geq\left|\lambda_{2}\right|\geq\left|\lambda_{3}\right|$; be $\mathbf{v_{1}}$ a unitary eigenvector associated with $\lambda_{1}$ at $\mathbf{s}$, it can be shown that $\mathbf{v_{1}}=\pm\mathbf{n}_{\mathcal{S}}$ for every $\mathbf{s}\in\mathcal{S}$. Therefore, we further condition the solution to Eq. 2 by adding an extra boundary condition enforcing alignment between the unitary directions associated to the maximum curvature field and the unitary normal field of the surface.
@@ 3.3 Implicit function learning
Similarly to previous works [26, 7], we extend this conditioning to points far from the surface ${\mathcal{S}}$. This is achieved by computing an approximation of the function $\mathit{t_{S}}$ (Eq. 1) based on nearest neighbors. Neumann’s boundary condition is expressed in the loss term:
@@ 3.4 Isosurface refinement
where $\mu(\cdot)$ is the mean value and $\sigma(\cdot)$ the standard deviation.
By minimizing the mean and variance of the learned function values at the isosurface, we ensure they are as close to zero as possible. This refinement process effectively reduces the oscillations and deviations that can occur at the critical boundary, thereby enhancing the overall quality of the reconstructed surfaces.
@@ 4.2 Network architecture
Our network takes spatial coordinates $x,y,z$ as input, and outputs the predicted $t_{\mathcal{S}}$ value from Eq. 1. The architecture consists of an 8-layer 256-units MLP with sine activations. The same number of layers and parameter count were used for baseline methods. Loss weights were experimentally set to $\lambda_{e},\lambda_{d},\lambda_{n}=1e^{4}$, $\lambda_{g}=1e^{3}$ and $\lambda_{\mu},\lambda_{\sigma}=1e^{5}$. Models were implemented in Python using Pytorch library, and distance metrics were computed using Pytorch3D [31]. We trained each network for 3,000 iterations, the first third of them utilized a learning rate of $1\mathrm{e}^{-4}$, in the following third we lowered it to $1\mathrm{e}^{-5}$, and in the final iterations (only the refinement loss active) we set the learning rate to $1\mathrm{e}^{-7}$ with a cosine decay. These final small learning rates allowed the refinement of the isosurface without losing the gained accuracy in the rest of the field. Regarding the value of $\alpha$ (Eq. 1), we experimentally found that choosing a value of 100 kept a good balance between training accuracy and reconstruction quality. Experiments were run on Ubuntu, an Intel(R) Xeon(R) Silver 4310 CPU, 256Gb of RAM, and 2 Nvidia A100 graphics cards of 80Gb VRAM.
@@ 4.3 Mesh reconstruction
For a thorough comparison of our method against existing techniques, we undertake mesh reconstruction using two gradient-based Marching Cubes algorithms, referred here as MC1 [46] and MC2 [14]. These methods perform linear interpolation between grid values for accurate placement of triangle vertices. Given this requirement and considering that the function $t_{\mathcal{S}}$, as described in Equation 1, exhibits quadratic behavior near the isosurface, we retrieve the true unsigned distance by first dividing by $\alpha$ and then applying the square root. For a comprehensive analysis we include SIREN [35], which does not approximate an unsigned distance, as illustrated in Figure 1. In this case, we used traditional Marching Cubes[19] for the reconstruction.
@@ 4.4.1 Closed shapes
We benchmarked DUDF against state-of-the-art methods for representing closed surfaces, where signed distances are well-defined. Given that our primary focus is not on closed surface representation, we limited this comparison to the ShapeNet car dataset [3], modifying the meshes for closure and omitting internal structures [40]. The comparative results are detailed in Table 1. While DeepSDF [27] excels in speed, it falls short in reconstruction quality. Being primarily designed for shape generation, DeepSDF is outperformed by methods tailored for accurate single shape representation, as evidenced by the Chamfer distance metrics between ground truth and reconstructed point clouds.
@@
SIREN [35] offers improved shape representation with enough detail and smoothness. However, since its zero level set satisfies the implicit function theorem, it cannot properly represent shapes with right angles and thin structures. In contrast, CAP-UDF [46] can better approximate these sharp features because it does not rely on continuously differentiable implicit neural representation networks, thereby achieving higher accuracy. However, the training time is high and the reliance on gradient based Marching Cubes methods for surface reconstruction tends to create surfaces with small holes and discontinuities which hinders the visual quality in some cases. Our approach captures a differentiable function’s zero level set like SIREN, but uniquely bypasses the implicit function theorem by having null gradients at the isosurface. This allows our method to effectively represent sharp features and complex geometries, leading to improved precision in surface reconstruction, particularly in scenarios where surfaces are not inherently smooth.
@@ 4.4.2 Open surfaces
When evaluating open-surface representation, we primarily compare our method against CAP-UDF [46]. This approach was shown to be more effective than earlier techniques (i.e. NDF [6]), that required a complex process for creating a dense point cloud and reconstructing a triangulated surface using the Ball Pivoting algorithm [1], which lead to poor quality meshes with great amounts of holes and irregularities Additionally, a recent method by Zhou et al. [44] has introduced constraints on the zero level set, which the authors claim it leads to better reconstructions. However, the absence of available implementation code in their official repository precluded a fair comparison with our method. Therefore, we concentrate our analysis on the most recent and available methodologies.
@@
time(s) $\downarrow$ L1CD $\downarrow$ L2CD $\downarrow$
$376$ $27.3$ $1.980$
$1390$ $18.1$ $1.110$
$\mathbf{326}$ $\mathbf{9.01}$ $\mathbf{0.025}$
$9.14$ $0.027$
@@
time(s) $\downarrow$ L1CD $\downarrow$ L2CD $\downarrow$
$374$ $40.5$ $8.810$
$1440$ $18.5$ $1.190$
$\mathbf{318}$ $\mathbf{8.70}$ $\mathbf{0.024}$
$8.82$ $0.026$
@@
time(s) $\downarrow$ L1CD $\downarrow$ L2CD $\downarrow$
$751$ $16.9$ $0.170$
$1040$ $\mathbf{10.9}$ $0.071$
$\mathbf{317}$ $12.3$ $\mathbf{0.057}$
$13.7$ $0.081$
Table 2: Training time, L1, and L2 mean Chamfer distances (×$\times 10^{3}$) for the evaluated open surface data sets. SIREN was trained as described in the original paper, while CAP-UDF and our method were trained on unsigned distances.
Qualitative comparisons can be observed in Fig. 3 and a full quantitative analysis is presented in Table 2. On the one hand, our method demonstrates a significant improvement in efficiency, consuming up to an order of magnitude less computational time than CAP-UDF, while also showing enhanced performance across all three data sets. This advantage is particularly evident in the DeepFashion and Multi-Garment data sets, where CAP-UDF tends to inaccurately close openings (like those at the ends of sleeves in clothing). On the other hand, since SIREN was devised for closed surfaces, it creates a negative distance valued shell around the isosurface. This causes to either close every hole or a thickened surface enclosing the ground-truth level set.
In addition to these experiments, we compare our method with CAP-UDF in the context of sphere tracing rendering. This comparison is presented in Fig. 4, where we showcase two rendering examples. Since learned unsigned distance fields in CAP-UDF do not grow linearly away from the surface, the marching steps in sphere tracing often fail to accurately intersect the surface, leading to undesirable visual artifacts. Even when hitting the surface, gradients might be undefined, leading to noisy shaded images. In contrast, our proposed distance function does not exhibit these problems, offering a more robust framework for direct rendering. Furthermore, our ability to compute normals using the maximum curvature field is useful for shading purposes, enhancing the rendering quality without undergoing an intermediate 3D mesh reconstruction step.

### v2

@@ 1 Introduction
Surface representation is a fundamental aspect in the field of 3D geometry processing, with explicit methods such as meshes, point clouds, and voxelized representations being traditional choices. Implicit surface representations, on the other hand, have been an integral part of the graphics pipeline for many decades. They encapsulate surfaces as the zero-level set of a function, providing a compact and continuous geometry representation. The novelty in recent years has emerged from parameterizing these implicit functions with Neural Networks (NNs), combining their learning capabilities with the advantages of implicit representations.
@@
We evaluate our framework with experiments that rigorously assess performance on a variety of challenging data sets and compare to competitive benchmarks in the field. Our method overcomes several limitations of previous approaches that struggled with the non-differentiability of UDFs. Results demonstrate that our approach not only enhances the reconstruction quality of open surfaces but also accelerates the training process. We believe the dual improvement in precision and efficiency enhance the applicability of neural UDFs representations in geometric processing for real-world scenarios.
@@ 2 Related Works
Closed surfaces. Recent years have seen a surge in the use of Neural Implicit Functions (NIFs) for modeling and reconstructing 3D shapes [39]. These representations are usually implemented with Multi-Layer Perceptrons (MLP) or Convolutional Neural Networks (CNNs), but differ in the learning task [21, 4, 29, 5]. On the one hand, some methods [21, 4, 29, 22, 5] learn an indicator function or binary occupancies which are used to reconstruct the 3D surface. On the other hand, alternative methods focus on estimating the SDF at any point in 3D space [27, 15, 23, 9, 11, 24]. The landscape of NIFs has further expanded with the introduction of novel approaches like implicit moving least-squares surfaces [16], a differentiable Poisson solver [28], a complex Gabor wavelet [32], and a level set alignment loss [20]. Notably, central works in the field [13, 35] approach the problem as finding solutions to the Eikonal equation alongside expressive boundary conditions, building upon the literature on solving PDEs with neural networks [34, 30]. In particular, SIREN [35] achieves remarkable results utilizing periodic activation functions, allowing to successfully control the function’s differential fields. Despite these advancements, a fundamental limitation of current NIF approaches remains in their inability to represent open surfaces, a characteristic often exhibited by real-world objects like scene walls, clothing, or vehicles with inner structures.
@@
Open surfaces. In order to model general non-watertight surfaces, Chibane et al. introduced the idea of learning UDFs, pioneering the handling of open surfaces with neural networks [6]. Following this, several methods have aimed to improve the performance of open-surface representation using neural networks [36, 37, 42, 41, 43, 45]. For instance, GIFS [42] and Neural Vector Fields [41] model the relationships between every two points instead of the relationships between points and surfaces. HSDF separately learns sign and distance fields to handle surfaces with arbitrary topologies[38]. NeuralUDF [18] and NeUDF [17] focus on learning an unsigned distance field by volume rendering for multi-view reconstruction of surfaces with arbitrary topologies. Closer to our work is CAP-UDF [45], that optimizes models on raw point clouds by learning to move 3D point queries until reaching the surface with a field consistency constraint. Despite these developments, since UDFs are not differentiable at the zero level set, the gradient field is ill-defined near the isosurface. This leads to difficulties during training, but specially during surface reconstruction, where gradients are often used to determine surface presence between points [14, 46], compute surface normals, and project points onto the isosurface [6] through gradient descent, among others. Recent work by Zhou et al. aimed to solve this issue by introducing constraints at the zero level set in the form of losses leading to a smoother surface [44]. In contrast, our approach formulates an Eikonal problem which is solved by learning a scaled distance function that remains differentiable close to the surface. We demonstrate that our method not only enhances reconstruction quality, addressing the prevailing challenges associated with the non-differentiability of the unsigned distance, but also improves state-of-the-art in terms of training time performance.
@@ 3.1 Mathematical background
Methods addressing closed surfaces approach the learning of SDFs as finding the solution to a system of Partial Differential Equations (PDEs) governed by the homogeneous Eikonal equation with Dirichlet and Neumann boundary conditions. Formally, given a closed surface $\mathcal{S}$ in $\mathcal{C}\subseteq\mathbb{R}^{3}$ they seek to find a continuous function $f$ that satisfies $\left\|\nabla f\right\|=1$, with boundary conditions $f_{\rvert_{\mathcal{S}}}=0$, $\nabla f_{\rvert_{\mathcal{S}}}=\mathbf{n}_{\mathcal{S}}$. Where $\mathbf{n}_{\mathcal{S}}$ denotes the unitary normal field at the surface. Although SDFs are not differentiable at every point and represent only a weak solution to the Eikonal equation, recent work has demonstrated considerable success in solving this problem with continuously differentiable implicit neural representation networks [35, 13, 26, 7]. These architectures use periodic activations, facilitating smoother optimization processes and improved control over the solution’s gradient field. This is achievable because locations where the signed distance lacks differentiability are distant from the isosurface, hence the approximation tends to be good in a close neighborhood of the zero level set. In particular, these networks enable accurate computation of critical topological properties, including mean and Gaussian curvature [26]. This contrasts with networks featuring piecewise linear activations, such as ReLUs, which have null second-order derivatives [35].
@@ 3.2 Problem statement
Our first insight is to redefine the unsigned distance field through the application of a hyperbolic scaling. In light of this understanding, we propose to learn the parameters $\theta$ of a neural network $f_{\theta}$ with periodic activation functions [35] to approximate the function:
@@
where $\mathit{d}_{\mathcal{S}}$ is the unsigned euclidean distance to surface $\mathcal{S}$; and $\alpha$ is a constant value. The function $\mathit{t}_{\mathcal{S}}$ is a differentiable approximation of $\mathit{d}_{\mathcal{S}}$, whose zero level set is surface $\mathcal{S}$. Fig. 2 (a) illustrates hyperbolic scaling’s quadratic smoothing near the isosurface and the linear behavior in the distance. The parameter $\alpha$ controls the distance at which the smoothing occurs.
Our second insight is that the distance scaling enables the application of continuously differentiable implicit neural representation networks to solve an Eikonal equation, while retaining the UDF’s open surface representation capabilities. For this task, we aim to address a heterogeneous Eikonal equation, for which we know $t_{\mathcal{S}}$ (Eq. 1) is a weak solution:
@@
The problem outlined in Eq. 2 does not incorporate the information of the surface normals. This contrasts with prior research focused on closed surfaces, where leveraging such information has led to improved accuracy and enhanced reconstruction quality by further constraining the set of feasible solutions. Our third key insight is that although the gradient of $t_{\mathcal{S}}$ (Eq. 1) becomes null at the isosurface, the direction in which the gradient norm increases most rapidly remains non-null. This direction is indicated by the maximum curvature of $f$, defined as the eigenvector associated with the largest eigenvalue of the Hessian matrix. Note that in this context, the maximum curvature does not refer to the surface’s curvature, but rather to the curvature of the hyperbolic scaled unsigned distance field $f$, where the function’s rate of change varies most rapidly. Formally, $\mathbf{H}_{t_{\mathcal{S}}}$ is the Hessian matrix of $t_{\mathcal{S}}$ which has 3 real eigenvalues $\left|\lambda_{1}\right|\geq\left|\lambda_{2}\right|\geq\left|\lambda_{3}\right|$; let $\mathbf{v_{1}}$ be a unitary eigenvector associated with $\lambda_{1}$ at $\mathbf{s}$, it can be shown that $\mathbf{v_{1}}=\pm\mathbf{n}_{\mathcal{S}}$ for every $\mathbf{s}\in\mathcal{S}$. Therefore, we further condition the solution to Eq. 2 by adding an extra boundary condition enforcing alignment between the unitary directions associated to the maximum curvature field and the unitary normal field of the surface.
@@ 3.3 Implicit function learning
Similarly to previous works [26, 7], we extend this conditioning to points far from the surface ${\mathcal{S}}$. This is achieved by computing an approximation of the function $t_{\mathcal{S}}$ (Eq. 1) based on nearest neighbors. Neumann’s boundary condition is expressed in the loss term:
@@ 3.4 Isosurface refinement
where $\mu(\cdot)$ is the mean value and $\sigma(\cdot)$ the standard deviation. By minimizing the mean and variance of the learned function values at the isosurface, we ensure they are as close to zero as possible. This refinement process effectively reduces the oscillations and deviations that can occur at the critical boundary, thereby enhancing the overall quality of the reconstructed surfaces.
@@ 4.2 Network architecture
Our network takes spatial coordinates $x,y,z$ as input, and outputs the predicted $t_{\mathcal{S}}$ value from Eq. 1. The architecture consists of an 8-layer 256-units MLP with sine activations. The same number of layers and parameter count were used for baseline methods. Loss weights were experimentally set to $\lambda_{e},\lambda_{d},\lambda_{n}=1e^{4}$, $\lambda_{g}=1e^{3}$ and $\lambda_{\mu},\lambda_{\sigma}=1e^{5}$. Models were implemented in Python using Pytorch library, and distance metrics were computed using Pytorch3D [31]. We trained each network for 3,000 iterations, the first third of them utilized a learning rate of $1\mathrm{e}^{-4}$, in the following third we lowered it to $1\mathrm{e}^{-5}$, and in the final iterations (only the refinement loss active) we set the learning rate to $1\mathrm{e}^{-7}$ with a cosine decay. These final small learning rates allowed the refinement of the isosurface without losing the gained accuracy in the rest of the field. Regarding $\alpha$ (Eq. 1), we experimentally found that choosing a value of 100 kept a good balance between training accuracy and reconstruction quality. Experiments were run on Ubuntu, an Intel(R) Xeon(R) Silver 4310 CPU, 256Gb of RAM, and 2 Nvidia A100 graphics cards of 80Gb VRAM.
@@ 4.3 Mesh reconstruction
For a thorough comparison of our method against existing techniques, we undertake mesh reconstruction using two gradient-based Marching Cubes algorithms, referred here as MC1 [46] and MC2 [14]. These methods perform linear interpolation between grid values for accurate placement of triangle vertices. Given this requirement and considering that the function $t_{\mathcal{S}}$, as described in Equation 1, exhibits quadratic behavior near the isosurface, we retrieve the true unsigned distance by first dividing by $\alpha$ and then applying the square root. We present results using the Chamfer distance (L1,L2), pervasive in related literature. Additionally, we compute the Normal Consistency metric (NC) as the mean value of the absolute cosine similarities between each point to its closest neighbor in the other point cloud.
@@ 4.4.1 Closed shapes
We benchmarked DUDF against state-of-the-art methods for representing closed surfaces, where signed distances are well-defined. Given that our primary focus is not on closed surface representation, we limited this comparison to the ShapeNet car dataset [3], modifying the meshes for closure and omitting internal structures [40]. The comparative results are detailed in Table 1.
While DeepSDF [27] excels in speed, it falls short in reconstruction quality. Being primarily designed for shape generation, DeepSDF is outperformed by methods tailored for accurate single shape representation.
@@
Despite SIREN [35] not reporting performance on learning distance fields, we included it in our comparison for the sake of fairness and comprehensiveness, given that our method’s backbone is built upon its architecture. When tested on close surfaces, SIREN shows improved shape representation with enough detail and smoothness. However, since its zero level set satisfies the implicit function theorem, it cannot properly represent shapes with right angles and thin structures. In contrast, CAP-UDF [46] can better approximate these sharp features because it does not rely on continuously differentiable NNs, thereby achieving higher accuracy. However, the training time is high and the reliance on gradient based Marching Cubes methods for surface reconstruction tends to create surfaces with small holes and discontinuities which hinders the visual quality in some cases. Our approach captures a differentiable function’s zero level set like SIREN, but uniquely bypasses the implicit function theorem by having null gradients at the isosurface. This allows our method to effectively represent sharp features and complex geometries, leading to improved precision in surface reconstruction, particularly in scenarios where surfaces are not inherently smooth.
@@ 4.4.2 Open surfaces
When evaluating open-surface representation, we primarily compare our method against CAP-UDF [46]. This approach was shown to be more effective than earlier techniques (i.e. NDF [6]), that required a complex process for creating a dense point cloud and reconstructing a triangulated surface using the Ball Pivoting algorithm [1]. This often leads to poor quality meshes with holes and irregularities. Additionally, a recent method by Zhou et al. [44] has introduced constraints on the zero level set, which the authors claim it leads to better reconstructions. However, the absence of available implementation code in their official repository precluded a fair comparison with our method. Therefore, we concentrate our analysis on the most recent and available methodologies.
@@
time(s) $\downarrow$ L1CD $\downarrow$ L2CD $\downarrow$ NC $\downarrow$
$376$ $27.3$ $1.980$ $0.107$
$1390$ $18.1$ $1.110$ $0.080$
$\mathbf{326}$ $\mathbf{9.01}$ $\mathbf{0.025}$ $0.024$
$9.14$ $0.027$ $\mathbf{0.020}$
@@
time(s) $\downarrow$ L1CD $\downarrow$ L2CD $\downarrow$ NC $\downarrow$
$374$ $40.5$ $8.810$ $0.094$
$1440$ $18.5$ $1.190$ $0.083$
$\mathbf{318}$ $\mathbf{8.70}$ $\mathbf{0.024}$ $0.026$
$8.82$ $0.026$ $\mathbf{0.021}$
@@
time(s) $\downarrow$ L1CD $\downarrow$ L2CD $\downarrow$ NC $\downarrow$
$751$ $16.9$ $0.170$ $\mathbf{0.240}$
$1040$ $\mathbf{10.9}$ $0.071$ $0.288$
$\mathbf{317}$ $12.3$ $\mathbf{0.057}$ $0.387$
$13.7$ $0.081$ $0.304$
Table 2: Training time, L1 and L2 mean Chamfer distances (×$\times 10^{3}$), and Normal Consistency (NC) for the evaluated open surface data sets. SIREN was trained as described in the original paper, while CAP-UDF and our method were trained on unsigned distances.
Qualitative comparisons can be observed in Fig. 3 and a full quantitative analysis is presented in Table 2. On the one hand, our method demonstrates a significant improvement in efficiency, consuming up to an order of magnitude less computational time than CAP-UDF, while also showing enhanced performance across all three data sets. This advantage is particularly evident in the DeepFashion and Multi-Garment data sets, where CAP-UDF tends to inaccurately close openings (like those at the ends of sleeves in clothing). On the other hand, SIREN creates a negative distance valued shell around the isosurface. This causes to either close every hole or a thickened surface enclosing the ground-truth level set.
In addition to these experiments, we compare our method with CAP-UDF in the context of sphere tracing rendering. This comparison is presented in Fig. 4, where we showcase two rendering examples. Since learned unsigned distance fields in CAP-UDF do not grow linearly away from the surface, the marching steps in sphere tracing often fail to accurately intersect the surface, leading to undesirable visual artifacts. Even when hitting the surface, gradients might be undefined, leading to noisy images. In contrast, our proposed distance function does not exhibit these problems, offering a more robust framework for direct rendering. Furthermore, our ability to compute normals using the maximum curvature field is useful for shading purposes, enhancing the rendering quality without undergoing an intermediate 3D mesh reconstruction step.